OpenStudy (anonymous):

Tamara is reviewing recent orders at her deli to determine which meats she should order. She found that of 1,000 orders, 450 customers ordered turkey, 375 customers ordered ham, and 250 customers ordered neither turkey nor ham. Based on these data, how many of the next 1,000 customers will order both turkey and ham? Show your work and use complete sentences.

OpenStudy (texaschic101):

1000 people total Turkey (450) Ham (375) Both (x) neither (250) (450 - x) + (x) + (375 - x) + 250 = 1000 -- combine like terms - x + 1075 = 1000 -- subtract 1075 from both sides -x = 1000 - 1075 -x = - 75 x = 75 only turkey -- 450 - x = 450 - 75 = 375 only ham -- 375 - x = 375 - 75 = 300 both -- 75 neither -- 250 total = (375 + 300 + 75 + 250) = 1000 people (correct) 75 of the next 1000 customers will order both turkey and ham